what is LEMIS theorm in vectors. give some examples

Grade:8

1 Answers

Komal
8 years ago
Lami's theorem in statics states that

if three coplanar forces are acting on a same point and keep it stationary, then it obeys the relation


[\frac{A}{\sin(\gamma)}=\frac{B}{\sin(\beta)}=\frac{C}{\sin(\alpha)}]

where A, B and C are the magnitude of forces acting at the point (say O), and the values of

α, βandγ are the angles directly opposite to the forces C, B and A respectively.